OpenStudy (anonymous):

What is the lateral area of a rectangular prism if the base edges are 3 centimeters and 4 centimeters and the height is 6 centimeters? Answer 84 cm2 92 cm2 150 cm2 220 cm2

OpenStudy (anonymous):

i think its B

OpenStudy (hoblos):

it must be 2(3)(4)+2(3)(6)+2(4)(6)=108 but it is not one of the options!!

OpenStudy (anonymous):

humm

OpenStudy (anonymous):

\(2 \times (3+4) \times 6 \)

OpenStudy (anonymous):

So that's 84 sq cm. Option \( \boxed{1} \)

OpenStudy (anonymous):

ok thank u both lots

OpenStudy (hoblos):

hmm it sounds right.. i thought the areas of the top and the bottum must count but it seems not

OpenStudy (anonymous):

That's total surface area @hoblos.

OpenStudy (hoblos):

but how?!! the area of the total surface is the area of the SIX rectangles that cover it by your formula you find the areas of only 4 of them!

OpenStudy (anonymous):

Yes top and bottom are not considered as lateral surface.

OpenStudy (hoblos):

0k now i got it... thanks

OpenStudy (anonymous):

Glad to help hoblos.

OpenStudy (anonymous):

A prism has a right triangle as its base with a hypotenuse that measures 5 meters and legs that measure 3 meters and 4 meters. What is the surface area of this prism if the height of the prism is 6 meters? Answer 78 m2 84 m2 96 m2 102 m2

OpenStudy (anonymous):

i got 96 for that one am i right

OpenStudy (anonymous):

New thread sunset :)

OpenStudy (anonymous):

ok i posted it again and if u go there i will give u a medal